OpenStudy (anonymous):

Show that if the magnitude of A-B = the magnitude of A+B then A and B are perpindicular

OpenStudy (anonymous):

\[\left| A-B \right| = \left| A+B \right|\] That is what the equation should actually look like. I am unsure of how to start. New to vector algebra

OpenStudy (jamesj):

Now square both sides and remember that \[ |v|^2 = v \cdot v \]

OpenStudy (anonymous):

So I get \[\left| A \right|^2-\left| B \right|^2 = \left| A \right|^2+\left| 2 AB \right| +\left| B \right|^2\] correct?

OpenStudy (jamesj):

No. It is not the case that \[ ******WRONG \ \ \ \ \ |A+B|^2 = |A|^2 +|2AB| + |B|^2 \] What is the case is that \[ |A+B|^2 = (A+B).(A+B) = A.A + A.B + B.A + B.B \] where . is the dot product Now expand out the left and side.

OpenStudy (anonymous):

which should become \[2ABcos\] sorry if that did not come through so it becomes \[\left| A \right|^2 + 2ABcos \theta+ \left| B \right|^2\]

OpenStudy (jamesj):

Ok. That will also work but is a bit more subtle. In any case, rewrite the left hand side.

OpenStudy (anonymous):

\[\left| A \right|^2 - \left| B \right|^2 = \left| A \right|^2 +2ABcos \theta + \left| B \right|^2\]

OpenStudy (jamesj):

No \[ |A-B|^2 = (A-B).(A-B) = |A|^2 - 2|A||B|\cos\theta + |B|^2 \]

OpenStudy (anonymous):

Sorry I just reworked it out and got that my bad

OpenStudy (jamesj):

Hence if \[ |A-B| = |A+B| \] then \[ |A|^2 - 2|A||B|\cos\theta + |B|^2 = |A|^2 + 2|A||B|\cos\theta + |B|^2 \] hence ... you take it from here

OpenStudy (anonymous):

It ultimately becomes \[\left| A \right|\left| B\right|\cos \theta = 0\] which means the angle between them is 90 or 180 degrees and therefore they are perpindicular Thanks!

OpenStudy (jamesj):

Almost, certainly it must be that \( \cos\theta = 0 \) which means they are perpendicular (but not 180 degrees, yes?)

OpenStudy (anonymous):

Sorry 270

OpenStudy (jamesj):

Right

OpenStudy (anonymous):

Do you think you could help me with a similar problem?

OpenStudy (jamesj):

Post it and if I can help I will.

OpenStudy (anonymous):

Prove the law of sines using the cross product. Here is what I have done so far...  From there I am stuck

OpenStudy (jamesj):

Post as a new question if you don't mind.